Suppose you are starting an office-cleaning service. You have spent $315 on equipment. To clean each office you use $4 worth of supplies. You charge $25 per office. How many offices must you clean to break even?
Answer by stanbon(75887)

You have spent $315 on equipment. To clean each office you use $4 worth of supplies. You charge $25 per office. How many offices must you clean to break even?
----------
Let "x" be the number of offices you clean.
EQUATION:
(gain per hour)*x=315
(25-4)x=315
21x=315
x=15 (# of rooms you need to clean to break even)
Cheers,
Stan H.